OpenStudy (calculusxy):

Mariana is 10 years old and has grown 7 inches in 2 years. When she was 8 years old, she was 52 inches tall and began recording her height. Mariana writes and graphs a function that models her height in inches over time in years. What is the initial value for this function? A, 59 B. 3.5 C. 7 D. 52 I think that the answer choice is B (or 3.5)

Mehek (mehek14):

I believe you are correct because when she was 8, she was 52 inches when she is 10, she is 59 \(\dfrac{59-52}{10-8}=\dfrac{7}{2}=3\)

OpenStudy (calculusxy):

You mean \[\large \frac{ 7 }{ 2 } = 3.5\] ?

Mehek (mehek14):

oh ye my bad

OpenStudy (calculusxy):

Ok thank you for your help!

Mehek (mehek14):

\(\dfrac{59-52}{10-8}=\dfrac{7}{2}=3.5\)

Mehek (mehek14):

yw

OpenStudy (mathmale):

What does 3.5 represent? Please use units of measurement as well as numbers.

OpenStudy (mathmale):

I'd suggest re-reading the original question. What result does it ask for? Is 3.5 that result?